Factorise: => k - 3 + k • O => k •i3 • (i) 12 - 7x + 1 (ii) 2 + 7x + 3 (iii) 6x2 + 6x - 6 (iv) 3 - x - 4

Factorise:

=> k - 3 + k • O => k •i3 •

(i) 12 - 7x + 1 (ii) 2 + 7x + 3
(iii) 6x2 + 6x - 6 (iv) 3 - x - 4

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) 12x2 - 7x + 1= 12r - 4x - 3% + 1 = 4x(3% -1) - 1(& - 1) = (4x-1X3x -1) (ii) 2x2 + 7.x + 3 = 2r + 6x + % + 3 = + 3) + 1(.% + 3) = (2r + Xx + 3). (iii) 6%2 + 5x - 6 = 6%2 + 9% - 4x - 6 =3%(2.x + 3) - 2(2.x + 3) =Cat - 2X2r + 3). (iv) 3x2 -% - 4 = 3x2 - 4x + 3% - 4 = x<& -4) + 1(& -4) = (x + 1X3x - 4).